A painting is 10 inches tall and 15 inches wide. A print of the painting is 25 inches tall. How wide is the painting?
Answer by checkley77(12844)

10/25=15/X
10X=25*15
10X=375
X=37.5 INCHES WIDE IS THE PRINT.